OpenStudy (anonymous):

Find the coordinates for the vertex of the graph of the function y= 3x^2-12x+12.

OpenStudy (anonymous):

f(x)=ax^2+bx+c vertex points (x1,y1) x1=-b/2a so 12/6=2 , y1=f(x1) so f(2)=3*2^2-12*2+12=0 vertex (2,0)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Thank you? He just quoted some random result and didn't show you why it is true :/ http://tiny.cc/po3k1

OpenStudy (anonymous):

:/